A long distance trucker travels 144 miles in one direction during a snow storm. Their return trip in rainy weather was completed at double speed and took 3 hours less time. What is the speed?
Answer by geetha_rama(94)

Ditance = 144 miles
time taken in one directtion = t hours
speed in one direction = s miles/hour
st = 144 -> equation 1
speed in return direction = 2s miles/hour
time taken in return direction = t-3 hours
2s(t-3) = 144 - > equation 2
from equation 1 & 2,
st = 2s(t-3)
t = 2t - 6
t = 6 hrs
s = 144/t
144/6
24 miles/hour
In first direction speed = 24 miles/hour and in return speed = 48 miles/hour